### v1

# AI Mirage: The Impostor Bias and the Deepfake Detection Challenge in the Era of Artificial Illusions

Abstract
This paper provides a comprehensive analysis of cognitive biases in forensics and digital forensics, examining their implications for decision-making processes in these fields. It explores the various types of cognitive biases that may arise during forensic investigations and digital forensic analyses, such as confirmation bias, expectation bias, overconfidence in errors, contextual bias, and attributional biases. It also evaluates existing methods and techniques used to mitigate cognitive biases in these contexts, assessing the effectiveness of interventions aimed at reducing biases and improving decision-making outcomes. Additionally, this paper introduces a new cognitive bias, called “impostor bias”, that may affect the use of generative Artificial Intelligence (AI) tools in forensics and digital forensics. The impostor bias is the tendency to doubt the authenticity or validity of the output generated by AI tools, such as deepfakes, in the form of audio, images, and videos. This bias may lead to erroneous judgments or false accusations, undermining the reliability and credibility of forensic evidence. The paper discusses the potential causes and consequences of the impostor bias, and suggests some strategies to prevent or counteract it. By addressing these topics, this paper seeks to offer valuable insights into understanding cognitive biases in forensic practices and provide recommendations for future research and practical applications to enhance the objectivity and validity of forensic investigations.
keywords:
Forensic Sciences, Cognitive Biases, Cognitive Psychology, Digital Forensics, Synthetic Data, Impostor Bias, Generative AI, GAN, Diffusion Models, Deepfake Detection
PACS:
$0$, $1$
MSC:
$0$, $1$
[…]
[inst1]organization=Department of Mathematics and Computer Science, University of Catania,addressline=Viale Andrea Doria 6, city=Catania, postcode=95126, state=CT, country=Italy
[…]
[inst2]organization=Department of Educational Sciences, Section of Psychology, University of Catania,addressline=Via Teatro Greco 84, city=Catania, postcode=95124, state=CT, country=Italy
1 Introduction
In the realm of forensic sciences, perhaps more than in other disciplines, the so-called objectivity of judgment in analyzing data and interpreting them for justice purposes is one of the crucial aspects. Beyond the specific technical aspects of individual disciplines, the authoritativeness of studies, and the competence of experts, there is a need to be informed about the so-called cognitive biases. Cognitive biases, rather than being deficits, are systematic preferences that influence the way we process, select, and retain information (Grisham et al., 2014; Lester et al., 2011). These biases can significantly impact our judgment and decision-making processes.
The tendency to jump to conclusions involves making hasty judgments or drawing premature conclusions without adequate evidence (Warman et al., 2013). Confirmation bias is another common cognitive bias where individuals selectively pay attention to information that aligns with their pre-existing beliefs, often disregarding contradictory evidence (Peters, 2020). Expectation bias, also known as experimenter’s bias, occurs when an individual’s anticipated outcome influences the actual outcome observed (Forensic Science Regulator, 2020). Overconfidence in errors is a related bias where individuals exhibit unwarranted certainty in their judgments and predictions, even when proven incorrect (Han, 2020; Sauvé et al., 2020). Contextual bias refers to situations where an individual’s judgment is influenced, either consciously or subconsciously, by additional information beyond what is being directly evaluated (Forensic Science Regulator, 2020). Attributional biases involve the unfair attribution of external events or experiences to oneself or others, thereby influencing our perception of causality (Sauvé et al., 2020). Cognitive biases can have both positive and negative effects, depending on the context and situation. They can facilitate swift decision-making when time is critical but can also lead to poor decisions and adverse outcomes (Berthet, 2022). Cognitive biases have been found to significantly influence decision-making processes in various professional fields such as management, finance, medicine, law, and psychology (Acciarini et al., 2021; Berthet, 2022; Neal and Grisso, 2014). Overconfidence bias, in particular, is a recurring bias that impacts decisions in these areas (Grežo, 2021; Wu et al., 2012). A summary of these biases can be found in Table 1.
Bias Description
Jump to Conclusions
Making quick judgments or drawing premature conclusions without sufficient evidence.
Confirmation Bias
Selectively paying attention to information that confirms our pre-existing beliefs while ignoring contradictory evidence.
Expectation Bias
An individual’s anticipated outcome influences the actual outcome that is observed.
Overconfidence in Errors
Individuals display an unwarranted certainty in their judgments and predictions, even when they are proven wrong.
Contextual Bias
An individual’s judgment is influenced (either consciously or subconsciously) by additional information beyond what is being directly evaluated.
Attributional Biases
Unfairly attributing external events or experiences to oneself or others, influencing our perception of causality.
Table 1: Summary of Cognitive Biases.
This paper delves into the influence of cognitive biases in the realm of forensics and digital forensics, with a particular focus on the detection of deepfakes. Deepfakes, which are Artificial Intelligence (AI)-generated images, videos, and audio, have a substantial impact on these fields. They can shape attitudes even when viewers are cognizant of the fact that they are viewing deepfakes, a phenomenon attributable to cognitive biases. Deepfakes pose threats such as the manipulation of public opinion and the impersonation of individuals, which could potentially lead to financial and reputational damage. In this context, we will introduce the concept of the impostor bias, which we define as an inherent distrust concerning the authenticity of multimedia elements like videos, photos, and audio. This distrust stems from the awareness that these elements can be realistically generated by AI models. Effective detection of deepfake products is crucial to prevent falling prey to the impostor bias and to avoid the potential confusion between real and fake multimedia products. Such confusion can lead to erroneous decisions and perceptions across all layers of multimedia evaluation in digital forensics. The challenges presented by deepfakes necessitate ongoing research and the development of advanced solutions. Consequently, we will analyze the most recent and efficient deepfake detection systems to aid operators and investigators in distinguishing between real and fake multimedia.
Particularly, the following points encapsulate the salient findings of this article:
  • 1.
    Cognitive biases in digital forensics: we discuss how cognitive biases can affect the perception and judgment of digital forensic investigators, especially in the face of complex and large-scale data.
  • 2.
    Deepfake detection methods: the state-of-the-art methods for detecting deepfakes, which are synthetic media created by advanced AI technologies, such as GANs and DMs, are reviewed.
  • 3.
    The impostor bias: we unveil the new concept of the impostor bias, which is the tendency to doubt the authenticity of real media due to the proliferation of deepfakes and the difficulty of distinguishing them from reality.
  • 4.
    Biases mitigating strategies: some strategies to reduce the impact of cognitive biases in digital forensics, such as using objective and standardized procedures, are proposed to enhance the training and education of forensic experts, and to adopt ethical and legal guidelines.
Finally, the paper is structured as follows: Section 1 introduced the concept of cognitive biases. Section 2 analyses their impact on forensic sciences. Section 3 explores some examples of cognitive biases in digital forensics, such as confirmation bias and pareidolia bias. Section 4 explores various bias mitigation strategies in both forensics and digital forensics. Section 5 presents the deepfakes and how they can be generated and managed. Section 6 introduces the impostor bias, a new type of bias triggered by AI media that affects the perception of reality. Section 7 reviews some of the most recent and relevant methods for deepfake detection, which is crucial to counter the impostor bias, as well as the problem of model attribution. Section 8 discusses the potential impact of impostor bias in digital forensics and everyday life. Section 9 concludes the paper and provides some directions for future research.
2 Cognitive Biases Impact on Forensic Sciences
Cognitive biases have been a concern in forensic science since 1984, when Larry Miller put forth a paper discussing the presence of bias in forensic document examiners (Stoel et al., 2014). He suggested the introduction of procedural modifications to reduce cognitive bias, which could potentially result in incorrect outcomes. These biases can significantly impact expert judgments and the criminal justice process (Stoel et al., 2014; Dror and Rosenthal, 2008; Neal and Grisso, 2014), are influenced by various factors, and can lead to errors and misinterpretation of evidence (Kassin et al., 2013; Cooper and Meterko, 2019; Bhadra, 2021). Both forensic experts and law enforcement professionals are susceptible to these biases, including confirmation bias (Thompson and Newman, 2015; Meterko and Cooper, 2022). The sources of bias can be categorized into three groups related to the case, the analyst, and human nature (Dror, 2020). Seven primary factors can introduce bias within the laboratory setting (Jeanguenat et al., 2017) (As synthetically sketched in Figure 1).
Figure 1: Seven potential sources of bias that could impact forensic decision-making.
Gardner et al. (2019) found that task-irrelevant information can bias forensic analysts’ decisions. Similarly, Nakhaeizadeh et al. (2014) discovered that irrelevant information can lead to confirmation bias in forensic anthropology. Dror et al. (2021) found that base-rate neglect could bias forensic pathologists’ decisions in child death cases. In DNA forensics, Jeanguenat et al. (2017) found that suspect-driven bias and rare alleles can influence interpretation. Neal et al. (2022) conducted a systematic review on cognitive biases and debiasing techniques in forensic mental health, finding significant bias effects. Lastly, Stevenage and Bennett (2017) found that irrelevant DNA test outcomes could bias fingerprint matching tasks, confirming the presence of contextual bias.
3 Exploring Cognitive Biases in Digital Forensics
The definition of Digital Forensic Science often referred to is the one from the Digital Forensic Research Workshop (DFRWS) in Palmer et al. (2001): “The use of scientifically derived and proven methods toward the preservation, collection, validation, identification, analysis, interpretation, documentation and presentation of digital evidence derived from digital sources for the purpose of facilitating or furthering the reconstruction of events found to be criminal, or helping to anticipate unauthorized actions shown to be disruptive to planned operations”. A key area within this field involves the proper acquisition of digital content such as images, videos, and audio to produce evidence for forensic investigations. Multimedia forensics focuses on verifying the authenticity of data and reconstructing the history of an image since its acquisition (Battiato et al., 2016; Fabio Arceri et al., 2023; Giudice et al., 2017; Piva, 2013). In the realm of digital forensics, cognitive biases can significantly impact the interpretation of data.
3.1 Confirmation Bias in Text and Face Recognition
Fontani’s blog post on Amped Software depicted a hypothetical situation featuring two characters, John and Lucy (Figure 2(Fontani, 2021). In this scenario, John inadvertently influenced Lucy’s interpretation of a license plate by prematurely sharing his own interpretation. This phenomenon, known as confirmation bias, may lead Lucy to unconsciously process pixels and select frames that align with John’s interpretation. The post further delved into the intricacies of face comparison, noting that varying processing techniques can result in significantly different facial appearances. It underscored the necessity of withholding the suspect’s face from the examiner prior to the enhancement process to prevent unconscious bias towards a match. Fontani also emphasized that different processing techniques can significantly alter facial appearances during enhancement (Figure 3). Therefore, it’s crucial that the examiner doesn’t see the suspect’s face before the enhancement process to avoid unconsciously adjusting the enhancement to create a match. This unconscious adjustment is a form of cognitive bias.
Figure 2: John asked Lucy for help with a license plate, reading “BC 537”, but unsure about the last two characters and the first one. Credit: Fontani (2021).
Figure 3: Image processing could lead to produce a noticeable different image, with different face characteristics. Credit: Fontani (2021).
Furthermore, Sunde and Dror (2019) underscored the importance of digital forensics as a rapidly growing field within forensic science. The authors analyzed seven specific sources of cognitive and human error within the digital forensics process and propose relevant countermeasures, concluding that while some cognitive and bias issues are common across forensic domains, others are unique and dependent on the specific characteristics of the domain, such as digital forensics.
3.2 Image Processing Could Lead to Pareidolia
A study by Di Lazzaro et al. (2013) focused on the potentially misleading effects of software techniques used for elaborating low-contrast images. The researchers used the Shroud of Turin, one of the most studied archeological objects in history, as an example (Figure 4). They demonstrated that image processing of both old and recent photographs of the Shroud could lead researchers to perceive inscriptions and patterns that do not actually exist. The study further emphasized that the limited static contrast of our eyes can make the perception of low-contrast images problematic. The brain’s ability to retrieve incomplete information can interpret false image pixels after image processing. This phenomenon, named “pareidolia”, can lead to the perception of patterns in Shroud photographs that do not exist in reality (Figure 5).
Figure 4: A supposed concealed image of a face on the back side of the Shroud is revealed through advanced image processing of a photograph published in a book. The image is flipped from right to left (b). A negative image of the face that can be seen on the front side of the Shroud, processed in the same way as (a). Credit: Fanti and Maggiolo (2004).
Figure 5: A magnified version of Figure 4b (a). A face resembling the Shroud that we discern in the top-left section of Figure 4b (as depicted on the right) (b). We can also discern another face in the bottom left section of Figure 4b. Pareidolia leads to false positives, enabling us to see faces in Figure 4b that aren’t actually there. Credit: Di Lazzaro et al. (2013).
The enhancement of images extracted from video cameras can lead to a degradation of the overall quality of information. This is due to factors such as excessive compression, distance from the recording plane, and limited overall resolution. International best practices suggest verifying on a case-by-case basis whether the level of information is sufficient to extract useful data for investigations (Khin et al., 2020). However, when examining low-contrast images that present pseudo-random visual patterns after an initial enhancement process, it is crucial to mitigate the risk of pareidolia. This bias is particularly potent when the object of interest refers to “human faces” or more generally to “letters/numbers” or known human structures (Wang and Yang, 2018; Zhou and Meng, 2020a). Pareidolia is a subconscious illusion that tends to associate random shapes with known forms, especially human figures and faces. Classic examples include seeing animals or human faces in clouds, or a human face on the moon. In forensic investigations, we also suggest to entrust the analysis and interpretation to automatic methods or experts who can follow a “blind testing” approach, i.e., an interpretation detached from the knowledge of details and the reference context. This helps to ensure the search for “certain” evidence is as rigorous and unbiased as possible.
Zhou and Meng (2020b) discussed how some individuals exhibit a looser decision criterion for detecting faces, making them more prone to perceive faces where none exist. This relates to a concept in signal detection theory known as response bias (Nguyen and Beins, 2013). In digital forensics, examiners may fall prey to response biases when analyzing ambiguous digital evidence, predisposing them to validate or dismiss forensic hypotheses based on non-diagnostic features. Just as some are more likely to see faces in random patterns due to biases in how they set thresholds for face judgments, forensic analysts could have biases influencing how strictly they apply standards of evidence to digital artifacts. Understanding individual differences in cognitive biases like threshold placement could help address potential sources of error and increase objectivity in digital forensic examinations.
Furthermore, we present a case study which involved the analysis of surveillance camera footage to determine the presence of a passenger in a vehicle involved in a legal case. The analysis included scrutiny of various cameras and raised concerns about the validity of conclusions drawn by experts. The frames are part of a detailed analysis of the vehicle’s passages from different surveillance cameras, considering image overlaps and real passage times, including sunset. The aim was to evaluate the methodologies and conclusions of technical consultants and expert witnesses regarding the analysis of vehicle images, highlighting the lack of scientific rigor and proposing new analyses suggesting the absence of a passenger in the vehicle. Specifically, Figure 6 shows a sequence of frames that have been processed by a technical consultant. The frames demonstrate a clear reflection effect that disappears as the vehicle moves. The effect is attributed to the ease with which contour lines can introduce aberrations. The frames also exhibit strong chromatic variability due to video compression, which is known to heavily penalize the color component.
Figure 6: Sequence of frames that have been processed by the technical consultant.
The technical advisor claimed to see the image of human faces in such images, a factor induced by confirmation bias: in reality, what is present in the images are pseudo-random blobs and their temporal hold is instead to be traced to a simpler and more obvious reflection. To demonstrate this, we conducted an experiment with college students. In the experiment, students were presented with frames from surveillance camera footage and tasked with identifying and interpreting the images. Crucially, no prior information was provided to the students to prevent any preconceived notions or confirmation biases from influencing their observations. The primary focus was on the students’ ability to discern whether a vehicle contained a passenger or not. The experiment was designed to rigorously assess the students’ cognitive perception, comprehension of the images, and their capacity to make accurate identifications solely based on the visual data at hand. This approach underscores the importance of maintaining an unbiased perspective in observational tasks. Table 2 summarizes the outcome of this analysis process, which reports a number of responses of 165 (15 subjects multiplied by the number of images viewed) in which in no case was the ”certain” presence of a human face evidenced.
Image Students’ Evaluation
Nothing: 8 Other: 7
Those who answered ”other” identified:
• Automobile elements
Nothing: 9 Other: 6
Those who answered ”other” identified:
• Automobile elements
Nothing: 10 Other: 5
Those who answered ”other” identified:
• Silhouette with sun reflection
• Automobile elements
Nothing: 9 Other: 6
Those who answered ”other” identified:
• A car seat and a person
• Reflected human silhouette
• A cow
Nothing: 13 Other: 2
Those who answered ”other” identified:
• Indistinct silhouettes behind glass
Nothing: 9 Other: 6
Those who answered ”other” identified:
• There are two human silhouettes
• Silhouettes of hands
Table 2: Table representing students’ evaluation of the various images cutted from Figure 6.
4 Bias Mitigating Strategies
Mitigating these cognitive biases could significantly enhance decision-making across society, promoting long-term human well-being. Extensive research has been conducted on whether and how these biases can be mitigated (Acciarini et al., 2021).
Korteling et al. (2021) conducted a systematic review of the literature on the retention and transfer of bias mitigation interventions, which yielded 12 peer-reviewed studies that adequately studied retention over a period of at least 14 days or transfer to different tasks and contexts. Most of these studies indicated that game-based interventions were effective after the retention interval and were more effective than video interventions. One study found indications of transfer of bias mitigation training across contexts.
4.1 Biases Mitigating Strategies in Forensics
Dror (2013) focused on strategies to mitigate these biases. He proposed that recognizing the spectrum of biases, not only those that can arise from knowing irrelevant case information, but also biases that emerge from base rate regularities, working “backwards” from the suspect to the evidence, and from the working environment itself, can strengthen forensic science. He suggested measures such as a triage approach and case managers to enhance the quality and effectiveness of the work carried out by forensic examiners.
In the pursuit of objectivity and accuracy in forensic analysis, the Forensic Science Regulator of the United Kingdom Government has proposed and implemented several strategies designed to ensure that the analysis is not influenced by any form of bias (Forensic Science Regulator, 2020). One of these strategies is the adoption of a structured and validated approach for the analysis and evaluation of evidence. Two models that exemplify this approach are the ACE-V model and the CAI model. The ACE-V model, which stands for Analysis, Comparison, Evaluation, and Verification, is a widely accepted method for comparing fingerprints (Reznicek et al., 2010). On the other hand, the CAI model, which stands for Case Assessment and Interpretation, is a framework based on Bayesian thinking (Jackson, 2011). It provides clarity on the role of forensic scientists within the criminal justice process and allows for the design of effective, efficient, and robust examination strategies. Documentation, independent verification, and maintaining ethical standards are other key strategies in the forensic process. Documentation involves taking contemporary notes and recording observations, decisions, and changes, ensuring transparency and enabling a thorough review. Independent verification of critical results by a competent peer, unaware of the original examiner’s judgment, adds an extra layer of scrutiny for accuracy. Furthermore, forensic operators must uphold ethical standards such as honesty, integrity, objectivity, and impartiality, and avoid conflicts of interest and external pressures to ensure unbiased and reliable analysis (Forensic Science Regulator, 2020). Lastly, controlling the flow of information to the analyst is paramount. This is achieved using methods such as “sequential unmasking” or “linear unmasking”, which help to prevent unnecessary information from influencing the analyst’s judgment. In this regard, Dror and Kukucka (2021) introduced Linear Sequential Unmasking–Expanded (LSU-E), a methodology aimed at improving decision-making in forensic science by reducing noise and bias. The approach advocates for an initial analysis of raw data or evidence without any reference material or context, followed by a sequential consideration of other relevant information based on its objectivity, relevance, and potential for bias. LSU-E is applicable to all types of decisions within forensic science, not just comparative ones. Its main objective is to optimize the sequence of information exposure to maximize information utility and minimize cognitive and psychological influences. In addition, LSU-E provides guidelines for determining the optimal sequence for presenting task-relevant information and emphasizes the need for transparent documentation of the influence and role of various pieces of information on the decision-making process.
Camilleri et al. (2019) proposed that while some laboratories have implemented bias mitigating strategies with varying impact on operational efficiency, there has been no systematic assessment of the risk posed by cognitive bias. They suggested using a risk management framework to assess the potential impact of bias on forensic interpretations across multiple disciplines. This approach proved useful in assessing the effectiveness of existing bias mitigating strategies and identifying the latent level of risk posed. Furthermore, the authors affirm that all forensic organizations should seek to implement bias limiting measures that are simple, cost-effective, and do not adversely impact efficiency. Recently, Douglass et al. (2023) provided an examination of cognitive biases in digital forensics, particularly focusing on the potential of video-recorded eyewitness identification procedures for assisting evaluators in assessing eyewitness accuracy. The authors conducted three studies involving a total of 1,630 participants. The results revealed that evaluators, on average, successfully differentiated accurate from inaccurate witnesses based on videos of identification procedures alone. However, this ability was compromised when extraneous incriminating evidence was also provided. The authors found that instructions highlighting the limitations of forensic evidence did not preserve evaluators’ ability to discern accuracy when extraneous incriminating case evidence was provided. Moreover, case information affected other judgments, such as perceptions of the witness’s view.
4.2 Biases Mitigating Strategies in Digital Forensics
Grubl and Lallie (2022) addressed one of the most significant challenges in digital forensics: the precise age estimation of victims in child sexual abuse and exploitation cases. Investigators often need to determine the age of victims by examining images and interpreting the sexual development stages and other human characteristics. However, this process can be negatively impacted by a large forensic backlog, cognitive bias, and the immense psychological stress that this work can entail. In response to these challenges, the authors evaluated existing facial image datasets and proposed a new dataset tailored to the needs of similar digital forensic research contributions. This dataset, which included 327 images of individuals aged 0 to 20, is tested on the Deep EXpectation algorithm pre-trained on the IMDB-WIKI dataset (Grubl and Lallie, 2022). The results for young adolescents aged 10 to 15 and older adolescents/adults aged 16 to 20 are very encouraging, achieving Mean Absolute Errors as low as 1.79.
Karie et al. (2019) also delved into the role of cognitive biases in digital forensics, particularly in the context of combating cybercrimes. The authors acknowledged that the increasing complexity of cyber-attacks and the vast amounts of data, also known as Big Data, that investigators need to sift through to unveil Potential Digital Evidence have made the process of forensic investigation more challenging. In response to these challenges, the authors proposed a generic framework for diverging DL cognitive computing techniques into Cyber Forensics (CF), referred to as the DLCF Framework. Such solutions can range from reducing bias in forensic investigations to challenging what evidence is considered admissible in a court of law or any civil hearing.
5 The Art and Science of Deepfakes
Figure 7: Statistics of papers published in the deepfake field. (a) Papers published from 2017 to 2023 with the keywords deepfake, deepfake creation, deepfake detection. (b) Numbers of papers published in: Article, Preprint, Proceedings, Chapter and Edited Book.
The advent of advanced AI technologies has introduced new challenges in this field. One such challenge is the detection of deepfakes, the research area of which is constantly expanding (as shown in Figure 7). Deepfakes are synthetic media created through generative models based mainly on Generative Adversarial Networks (GANs) and Diffusion Models (DMs) (Goodfellow et al., 2014; Sohl-Dickstein et al., 2015). GANs are composed of a Generator (G) and a Discriminator (D) trained simultaneously through a competitive process. The Generator is trained to capture the data distribution of the training set Ts. The Discriminator is trained to distinguish the images created by G from the set Ts. When G creates images with the same data distribution as Ts, D will no longer be able to solve its task and the training phase can be considered completed. Currently, researchers demonstrated that synthetic images created by DMs (Sohl-Dickstein et al., 2015) are better than those generated by GAN engines in terms of photorealism, as the creation process follows a more accurate and “controlled” flow. The basic idea of DMs is to iteratively add noise to an input random noise vector for synthetic data generation in order to model complex data distributions. Figure 8 and Figure 9 show generic GAN and DM schemes related to the creation of synthetic people’s faces.
Deepfakes can pose significant challenges in distinguishing real images from manipulated ones, thereby complicating the task of digital forensic investigators: in fact, the problem of deepfake detection has been addressed extensively by the scientific community (Masood et al., 2023; Verdoliva, 2020). In this context, preventing cognitive biases in digital forensics becomes crucial to ensure the objectivity and neutrality of judgments.
Figure 8: A standard GAN framework including a Generator (G) that creates data samples from random noise, aiming to mimic the training set’s data distribution; a Discriminator (D) differentiates between real and G-generated data. Both are trained competitively.
Figure 9: A generic Diffusion Model learning a latent variable model that maps the latent space with a fixed Markov chain. The data is corrupted by adding Gaussian noise gradually to get the approximate posterior $q(x_{t}|x_{t-1})$. The aim of training is to learn the inverse process $(p_{\theta}(x_{t-1}|x_{t}))$ so that new data can be generated by running the chain backward. The latent variables $x_{1}$, …, $x_{t}$ have the same dimension as the data $x_{0}$.
6 The Impostor Bias: How AI Media Triggers Bias and Doubt in Perception
The emergence of Generative AI (GenAI) has brought about a sea change in the multimedia landscape, opening up novel avenues for crafting and modifying content. It is largely attributed to the development of a class of machine learning models known as foundation models (FMs) (Rabowsky, 2023). These models, which include the likes of ChatGPT released in November 2022, marked the beginning of a new era in Artificial Intelligence. Foundation models are distinguished by their powerful applications to GenAI, which involves the use of models to generate new content and transform existing content (Rabowsky, 2023). GenAI models can produce high-quality artistic media for visual arts, concept art, music, fiction, literature, video, and animation; distinguishing the real from the fake is becoming increasingly complex, as in the case of human face recognition and its veracity (Figure 10).
Figure 10: Do these people exist? Credit: Guarnera et al. (2022)
The existence of GenAI and its knowledge can lead to the development of a new type of cognitive bias, which we have identified as the “impostor bias”. This refers to an a priori distrust regarding the veracity of multimedia elements such as videos, photos, and audio, due to the awareness that these can be realistically generated by Artificial Intelligence models. Being subject to this bias, a condition towards which we believe we are increasingly heading in parallel with the increase in realism of GenAI multimedia products, entails and will entail a constant doubt that will also be influential when the victims are, for example, court operators, investigators, or magistrates themselves. In this regard, the technologies previously analyzed regarding deepfake detection will increasingly have to orient themselves not only towards the recognition of manipulated or altered multimedia artifacts, but also those generated from scratch starting from descriptive textual inputs. This is a great challenge on which we will have to work to avoid falling into cognitive traps dictated by prejudices of technological suspicion, risking declaring or considering original multimedia products instead generated by AI, and vice versa. Another risk stemming from GenAI pertains to the counterfeiting of artworks, the sale of forged masterpieces, and the infringement of copyright laws. Indeed, such technologies are now capable of simulating the technique and artistic style of the most famous artists, thereby compromising the ability to correctly discern between real and simulated works (Epstein et al., 2023; Leotta et al., 2023) (Figure 11). Forensic examiners from 21 countries showed limited understanding and appreciation of cognitive bias, with fewer than half supporting blind testing, highlighting the need for procedural reforms to blind them to potentially biasing information (Kukucka et al., 2017).
Figure 11: Real vs deepfake images of famous artists.
7 Deepfake Detection Methods
Figure 12: Fourier spectrum of different categories of data: (a) real images; (b) images generated by two GAN architectures (GDWCT (Cho et al., 2019) and StarGAN (Choi et al., 2018)); (c) images generated by two DM architectures (Latent Diffusion (Ramesh et al., 2022) and GLIDE (Nichol et al., 2022)). The abnormal frequencies (light peaks) are mainly visible in the images generated by artificial intelligence engines.
Researchers have demonstrated that generative engines leave traces on synthetic content that can be identified and detected in the frequency domain (Guarnera et al., 2020c; Zhang et al., 2019; Marra et al., 2019; Giudice et al., 2021; Dzanic et al., 2020; Durall et al., 2020) (Figure 12). These traces are characterized by both the network architecture (number and type of layers) and its specific parameters (Yu et al., 2019a). In order to distinguish real data from deepfake, Guarnera et al. (2020a, b) proposed methods based on the Expectation-Maximization (Moon, 1996) algorithm capable of capturing traces defined as the correlation of pixels left by convolutional layers.
Wang et al. (2020) used ResNET-50 to distinguish real and ProGAN-generated images, showing generalization across different GANs. FakeSpotter, proposed by Wang et al. (2021), detects GAN-generated faces by monitoring CNN neuron behaviors. Vision Transformer-based solutions for deepfake detection have also been proposed (Wodajo and Atnafu, 2021; Coccomini et al., 2022; Heo et al., 2023; Wang et al., 2022)Wodajo and Atnafu (2021) combined transformers with a convolutional network to extract patches from detected faces in videos. Several studies (Yu et al., 2019b, 2021; Girish et al., 2021; Asnani et al., 2023; Yu et al., 2020; Guarnera et al., 2022) have explored identifying specific GAN models used in creation (Model Attribution Task). Guarnera et al. (2022) distinguished 100 StyleGAN2 instances using ResNET-18 (He et al., 2016) and metric learning (Liu et al., 2012), demonstrating the method’s effectiveness in deepfake model recognition.
The scientific community is also working extensively on the creation of advanced techniques for the detection of synthetic images created by diffusion models. Corvi et al. (2023) have been trying to understand how difficult it is to distinguish synthetic images generated from diffusion models from real ones, and whether current state-of-the-art detectors are suitable for this task. Sha et al. (2023) proposed DE-FAKE, a machine-learning classifier-based method for diffusion model detection on four popular text-image architectures. Guarnera et al. (2023) proposed a hierarchical approach based on different architectures in order to define: whether the image is real or manipulated via any generative architecture (AI-generated); the specific framework used among GAN or DM; defines the specific generative architecture used among a predefined set. Furthermore, another practical digital forensic tool, Transfer learning-based Autoencoder with Residuals (TAR), was proposed (Lee et al., 2021). The ultimate goal of TAR was to develop a unified model to detect various types of deepfake videos with high accuracy, with only a small number of training samples that can work well in real-world settings. A short summary of these methods can be found in Table 3.
Reference Generation Models Database(s) Used Precision (avg)
He et al. (2016) StyleGAN, StyleGAN2-ADA FFHQ (Flickr-Faces-HQ) 96.2%
Wang et al. (2020) ProGAN CelebA 99.1%
Guarnera et al. (2023)
GANs: (AttGAN, CycleGAN, GDWCT, IMLE, ProGAN,
StarGAN, StarGAN-v2, StyleGAN, StyleGAN2)
DMs: (DALL·E 2, GLIDE, Latent Diffusion, Stable Diffusion)
CelebA, FFHQ, ImageNet
97,6% (Level 1)
98,0% (Level 2)
97,8% (Level 3, GANs)
98,0% (Level 3, DMs)
Wang et al. (2021) StyleGAN, StyleGAN2, BigGAN, ProGAN FaceForensics++ 90.6%
Wodajo and Atnafu (2021)
FaceSwap, Face2Face, FaceShifter, NeuralTextures,
DeepFakeDetection
FaceForensics++, UADFV 91.5%
Lee et al. (2021) FaceSwap, Face2Face, DeepFake, NeuralTextures FaceForensics++
98.0% deepfake type detection
89.5% on DW videos
Sha et al. (2023) GLIDE, Latent Diffusion, Stable Diffusion, DALL·E 2 MSCOCO (a), Flickr30k (b) 90.2%(a), 84.6% (b)
Table 3: A summary of the discussed deepfake detection methods. Legend: AUC = Area Under the Curve; TAR = Transfer learning-based Autoencoder with Residuals; ViT = Vision Transformer.
Experimental results of all these methods show that, in general, all generative models leave unique traces that can solve all previously listed tasks with high accuracy. Therefore, these methods can be used in order to help the general user in countering what we called the impostor bias phenomenon. However, we want to highlight an important element of the previously listed methods and not just deepfake detection. All methods in the literature achieve extremely high results in ”constrained” contexts, that is, with architectures known a priori. In practice, current deepfake detectors fail to generalize with synthetic images generated by novel architectures (different from those used during the training procedure), resulting in a drastic drop in classification performance. Some recent new methods (Dong et al., 2023; Coccomini et al., 2023) published by the scientific community seem to be good starting points in order to achieve generalization.
8 Discussion
We highlighted the importance of deepfakes in the field of digital forensics, which we consider to be a topic of great relevance and timeliness: in particular, the concept of impostor bias that we introduced is increasingly prevalent, although it had not yet been named and defined precisely. A concrete example of impostor bias was seen with the recent war in Ukraine, the first war characterized by constant propaganda from both sides carried out on all media, and especially on social networks (Ciuriak, 2022; Suciu, 2022). Being flooded with videos, photos and statements inevitably leads to believe a priori that everything that is seen can be modified, or generated by AI models through a simple prompt (Linehan et al., 2023). The habit of suspicion developed during this “digital warfare” with propaganda and false information led to immediate doubt of all published media, sometimes leading to the discovery that some of this media was deepfake (Bond, 2023).
This condition of prior doubt on the veracity of the media will be more and more invasive and constant, especially when it comes to delicate material, such as a communication from a head of state in wartime (Linehan et al., 2023): for this reason, we believe that it is necessary to use the best tools for detecting deepfakes and AI-generated multimedia to avoid being deceived, especially to avoid considering fake multimedia that, paradoxically, are instead real.
Therefore, the data presented herein, along with previous findings, underscore the imperative need for further exploration of cognitive biases in digital forensic science. The focus should not be solely on the biases themselves, but rather on the strategies that can be employed to counteract and disengage from them. One potential approach could involve the creation and implementation of specialized training programs for forensic practitioners. These programs would aim to address cognitive biases in various scenarios, such as those examined in this review. By enhancing awareness of cognitive biases, we could potentially reduce the likelihood of practitioners being influenced or overwhelmed by these biases. This would enable them to disregard irrelevant elements of the investigation, separate personal experiences and other contextual factors that may “contaminate” decision-making processes, thereby preserving the objectivity of their judgments.
9 Conclusions and Future Works
This article explores cognitive biases in forensics and digital forensics, including confirmation bias, expectation bias, and overconfidence in errors. It discusses methods to reduce these biases, such as game-based interventions and the Linear Sequential Unmasking-Expanded approach. The article highlights the challenge of deepfakes, synthetic media that can manipulate or impersonate individuals, threatening the integrity of digital evidence. It surveys current deepfake detection methods and their limitations. The article introduces the impostor bias, a new cognitive bias that may lead to false negatives and reduced confidence in digital forensic findings due to the prevalence of deepfakes. We conclude by proposing new research directions in digital forensics, suggesting the creation of advanced deepfake detection methods, exploring the factors contributing to impostor bias, and recommending interventions to mitigate this bias. We also underscore the need to examine the ethical, legal, and social implications of deepfakes.
10 Acknowledgments
This research is supported by Azione IV.4 - “Dottorati e contratti di ricerca su tematiche dell’innovazione” del nuovo Asse IV del PON Ricerca e Innovazione 2014-2020 “Istruzione e ricerca per il recupero - REACT-EU”- CUP: E65F21002580005.

### v2

@@
-AI Mirage: The Impostor Bias and the Deepfake Detection Challenge in the Era of Artificial Illusions
+GenAI Mirage: The Impostor Bias and the Deepfake Detection Challenge in the Era of Artificial Illusions
 Abstract
-This paper provides a comprehensive analysis of cognitive biases in forensics and digital forensics, examining their implications for decision-making processes in these fields.
-It explores the various types of cognitive biases that may arise during forensic investigations and digital forensic analyses, such as confirmation bias, expectation bias, overconfidence in errors, contextual bias, and attributional biases.
-It also evaluates existing methods and techniques used to mitigate cognitive biases in these contexts, assessing the effectiveness of interventions aimed at reducing biases and improving decision-making outcomes.
-Additionally, this paper introduces a new cognitive bias, called “impostor bias”, that may affect the use of generative Artificial Intelligence (AI) tools in forensics and digital forensics.
-The impostor bias is the tendency to doubt the authenticity or validity of the output generated by AI tools, such as deepfakes, in the form of audio, images, and videos.
-This bias may lead to erroneous judgments or false accusations, undermining the reliability and credibility of forensic evidence.
-The paper discusses the potential causes and consequences of the impostor bias, and suggests some strategies to prevent or counteract it.
-By addressing these topics, this paper seeks to offer valuable insights into understanding cognitive biases in forensic practices and provide recommendations for future research and practical applications to enhance the objectivity and validity of forensic investigations.
+This paper examines the impact of cognitive biases on decision-making in forensics and digital forensics, exploring biases such as confirmation bias, anchoring bias, and hindsight bias.
+It assesses existing methods to mitigate biases and improve decision-making, introducing the novel “Impostor Bias”, which arises as a systematic tendency to question the authenticity of multimedia content, such as audio, images, and videos, often assuming they are generated by AI tools.
+This bias goes beyond evaluators’ knowledge levels, as it can lead to erroneous judgments and false accusations, undermining the reliability and credibility of forensic evidence.
+Impostor Bias stems from an a priori assumption rather than an objective content assessment, and its impact is expected to grow with the increasing realism of AI-generated multimedia products.
+The paper discusses the potential causes and consequences of Impostor Bias, suggesting strategies for prevention and counteraction.
+By addressing these topics, this paper aims to provide valuable insights, enhance the objectivity and validity of forensic investigations, and offer recommendations for future research and practical applications to ensure the integrity and reliability of forensic practices.
 keywords:
@@
 $0$ , $1$
+journal: Forensic Science International: Digital Investigation
 […]
@@
 1 Introduction
-In the realm of forensic sciences, perhaps more than in other disciplines, the so-called objectivity of judgment in analyzing data and interpreting them for justice purposes is one of the crucial aspects.
-Beyond the specific technical aspects of individual disciplines, the authoritativeness of studies, and the competence of experts, there is a need to be informed about the so-called cognitive biases.
-Cognitive biases, rather than being deficits, are systematic preferences that influence the way we process, select, and retain information (Grisham et al., 2014; Lester et al., 2011).
-These biases can significantly impact our judgment and decision-making processes.
-The tendency to jump to conclusions involves making hasty judgments or drawing premature conclusions without adequate evidence (Warman et al., 2013).
-Confirmation bias is another common cognitive bias where individuals selectively pay attention to information that aligns with their pre-existing beliefs, often disregarding contradictory evidence (Peters, 2020).
-Expectation bias, also known as experimenter’s bias, occurs when an individual’s anticipated outcome influences the actual outcome observed (Forensic Science Regulator, 2020).
-Overconfidence in errors is a related bias where individuals exhibit unwarranted certainty in their judgments and predictions, even when proven incorrect (Han, 2020; Sauvé et al., 2020).
-Contextual bias refers to situations where an individual’s judgment is influenced, either consciously or subconsciously, by additional information beyond what is being directly evaluated (Forensic Science Regulator, 2020).
-Attributional biases involve the unfair attribution of external events or experiences to oneself or others, thereby influencing our perception of causality (Sauvé et al., 2020).
-Cognitive biases can have both positive and negative effects, depending on the context and situation.
-They can facilitate swift decision-making when time is critical but can also lead to poor decisions and adverse outcomes (Berthet, 2022).
-Cognitive biases have been found to significantly influence decision-making processes in various professional fields such as management, finance, medicine, law, and psychology (Acciarini et al., 2021; Berthet, 2022; Neal and Grisso, 2014).
-Overconfidence bias, in particular, is a recurring bias that impacts decisions in these areas (Grežo, 2021; Wu et al., 2012).
-A summary of these biases can be found in Table 1.
-Bias Description
-Jump to Conclusions
-Making quick judgments or drawing premature conclusions without sufficient evidence.
-Confirmation Bias
-Selectively paying attention to information that confirms our pre-existing beliefs while ignoring contradictory evidence.
-Expectation Bias
-An individual’s anticipated outcome influences the actual outcome that is observed.
-Overconfidence in Errors
-Individuals display an unwarranted certainty in their judgments and predictions, even when they are proven wrong.
-Contextual Bias
-An individual’s judgment is influenced (either consciously or subconsciously) by additional information beyond what is being directly evaluated.
-Attributional Biases
-Unfairly attributing external events or experiences to oneself or others, influencing our perception of causality.
-Table 1: Summary of Cognitive Biases.
-This paper delves into the influence of cognitive biases in the realm of forensics and digital forensics, with a particular focus on the detection of deepfakes.
-Deepfakes, which are Artificial Intelligence (AI)-generated images, videos, and audio, have a substantial impact on these fields.
-They can shape attitudes even when viewers are cognizant of the fact that they are viewing deepfakes, a phenomenon attributable to cognitive biases.
-Deepfakes pose threats such as the manipulation of public opinion and the impersonation of individuals, which could potentially lead to financial and reputational damage.
-In this context, we will introduce the concept of the impostor bias, which we define as an inherent distrust concerning the authenticity of multimedia elements like videos, photos, and audio.
-This distrust stems from the awareness that these elements can be realistically generated by AI models.
-Effective detection of deepfake products is crucial to prevent falling prey to the impostor bias and to avoid the potential confusion between real and fake multimedia products.
-Such confusion can lead to erroneous decisions and perceptions across all layers of multimedia evaluation in digital forensics.
-The challenges presented by deepfakes necessitate ongoing research and the development of advanced solutions.
-Consequently, we will analyze the most recent and efficient deepfake detection systems to aid operators and investigators in distinguishing between real and fake multimedia.
+In forensic sciences, the objectivity of judgment in analyzing data for justice purposes is paramount.
+Beyond technical expertise, awareness of cognitive biases is crucial.
+These biases, rather than being deficits, are systematic preferences that influence the way we process, select, and retain information (Lester et al., 2011; Grisham et al., 2014).
+They can have both positive and negative effects, depending on the context and situation, facilitating swift decision-making when time is critical but also leading to poor decisions and adverse outcomes (Meterko and Cooper, 2022; Berthet, 2022).
+Some of these biases have been identified as a significant factor impacting the objectivity and accuracy of forensic science (Bhadra, 2021).
+For example, law enforcement professionals showed vulnerable to confirmation bias (Meterko and Cooper, 2022), which is a tendency to search for, interpret, and remember information in a way that confirms one’s preexisting beliefs or hypotheses (Nickerson, 1998).
+Another recurring bias in forensics is the anchoring bias or effect (Edmond et al., 2015), which occurs when an individual relies too heavily on an initial piece of information (the “anchor”) when making decisions (Chapman and Johnson, 1994, 1999).
+One more relevant bias related to forensic science is the hindsight bias (Giroux et al., 2016), which occurs when people believe that an event is more predictable after it becomes known, involving memory distortion, beliefs about objective likelihoods, and subjective beliefs about one’s own prediction abilities (Roese and Vohs, 2012).
+The discipline of digital forensics is not exempt from these challenges (Sunde and Dror, 2019).
+Consequently, an increasing number of scholarly investigations are focusing on this particular field (Sunde and Dror, 2021).
+This paper explores the impact of cognitive biases in forensics and digital forensics, with a focus on deepfakes and Artificial Intelligence (AI)-generated multimedia content, which pose threats such as manipulating public opinion and impersonating individuals.
+We introduce the Impostor Bias, an inherent distrust of multimedia authenticity due to the prevalence of AI-generated content.
+Effective detection of deepfakes is crucial to prevent confusion between real and fake multimedia, which can lead to erroneous decisions.
+We analyze advanced deepfake detection systems to aid operators in distinguishing authentic content, addressing the challenges posed by deepfakes, and ensuring accurate multimedia evaluation in digital forensics.
 Particularly, the following points encapsulate the salient findings of this article:
@@
 2 Cognitive Biases Impact on Forensic Sciences
-Cognitive biases have been a concern in forensic science since 1984, when Larry Miller put forth a paper discussing the presence of bias in forensic document examiners (Stoel et al., 2014).
+Cognitive biases have been a concern in forensic science since 1984, when Larry Miller published a work discussing the presence of bias in forensic document examiners (Stoel et al., 2014).
 He suggested the introduction of procedural modifications to reduce cognitive bias, which could potentially result in incorrect outcomes.
 These biases can significantly impact expert judgments and the criminal justice process (Stoel et al., 2014; Dror and Rosenthal, 2008; Neal and Grisso, 2014), are influenced by various factors, and can lead to errors and misinterpretation of evidence (Kassin et al., 2013; Cooper and Meterko, 2019; Bhadra, 2021).
-Both forensic experts and law enforcement professionals are susceptible to these biases, including confirmation bias (Thompson and Newman, 2015; Meterko and Cooper, 2022).
-The sources of bias can be categorized into three groups related to the case, the analyst, and human nature (Dror, 2020).
-Seven primary factors can introduce bias within the laboratory setting (Jeanguenat et al., 2017) (As synthetically sketched in Figure 1).
-Figure 1: Seven potential sources of bias that could impact forensic decision-making.
-Gardner et al.
-(2019) found that task-irrelevant information can bias forensic analysts’ decisions.
-Similarly, Nakhaeizadeh et al.
-(2014) discovered that irrelevant information can lead to confirmation bias in forensic anthropology.
+Both forensic experts and law enforcement professionals are susceptible to these biases, and one of the most common of them is the confirmation bias (Moser, 2013; Thompson and Newman, 2015; van den Eeden et al., 2019; Cooper and Meterko, 2019; Meterko and Cooper, 2022): if a forensic analyst believes a suspect is guilty, they might unconsciously interpret ambiguous evidence as incriminating (Kassin et al., 2013; van den Eeden et al., 2019)Gardner et al.
+(2019) found that task-irrelevant information can bias forensic analysts’ decisions, suggesting that extraneous details can inadvertently sway the interpretation of evidence.
+This is echoed by  Nakhaeizadeh et al.
+(2014), who discovered that irrelevant information can lead to confirmation bias in forensic anthropology, potentially leading to skewed conclusions based on pre-existing beliefs rather than objective evidence.
 Dror et al.
-(2021) found that base-rate neglect could bias forensic pathologists’ decisions in child death cases.
-In DNA forensics, Jeanguenat et al.
-(2017) found that suspect-driven bias and rare alleles can influence interpretation.
-Neal et al.
-(2022) conducted a systematic review on cognitive biases and debiasing techniques in forensic mental health, finding significant bias effects.
+(2021) further explored this issue, finding that base-rate neglect could bias forensic pathologists’ decisions in child death cases.
+This suggests that statistical information about the prevalence of certain causes of death may be overlooked, leading to potential misinterpretations.
+In the realm of DNA forensics, Jeanguenat et al.
+(2017) found that suspect-driven bias and the presence of rare alleles can influence interpretation.
+This highlights the potential for preconceived notions about a suspect, as well as the rarity of certain genetic markers, to affect the analysis of DNA evidence.
+In a study by Douglass et al.
+(2023), evaluators successfully differentiated accurate from inaccurate witnesses based on videos of identification procedures alone, but their ability to discern accuracy was disrupted when extraneous incriminating evidence was also provided.
+This aligns with confirmation bias, where evaluators tend to favor information that confirms their existing beliefs or expectations, even when it contradicts objective evidence.
+Regarding the anchoring bias, in a forensic context, an analyst might give undue weight to the first piece of evidence they examine, which could skew their interpretation of subsequent evidence (Edmond et al., 2015; Meterko and Cooper, 2022).
+For the hindsight bias, this could lead to overconfidence in the accuracy of a forensic analysis after a suspect has been identified (Giroux et al., 2016; Beltrani et al., 2018; Meterko and Cooper, 2022)Neal et al.
+(2022) conducted a systematic review on cognitive biases and debiasing techniques in forensic mental health, finding significant bias effects specifically for confirmation and hindsight bias.
 Lastly, Stevenage and Bennett (2017) found that irrelevant DNA test outcomes could bias fingerprint matching tasks, confirming the presence of contextual bias.
+This suggests that unrelated test results can influence the interpretation of fingerprint evidence.
+The sources of bias can be categorized into three groups related to the case, the analyst, and human nature.
+Inspired by the work of Dror et al.
+(Dror, 2020), we define a taxonomy of potential sources of bias.
+These biases can introduce cognitive distortions into the processes of sampling, observing, strategizing tests, analyzing, and drawing conclusions, even when conducted by experts.
+This taxonomy is synthetically sketched in Figure 1.
+Figure 1: Eight potential sources of bias that could impact forensic decision-making.
 3 Exploring Cognitive Biases in Digital Forensics
@@
 A key area within this field involves the proper acquisition of digital content such as images, videos, and audio to produce evidence for forensic investigations.
-Multimedia forensics focuses on verifying the authenticity of data and reconstructing the history of an image since its acquisition (Battiato et al., 2016; Fabio Arceri et al., 2023; Giudice et al., 2017; Piva, 2013).
-In the realm of digital forensics, cognitive biases can significantly impact the interpretation of data.
+Multimedia forensics focuses on verifying the authenticity of data and reconstructing the history of an image since its acquisition (Battiato et al., 2016; Arceri et al., 2023; Giudice et al., 2017; Piva, 2013).
+In the realm of digital forensics, cognitive bias emerges as a subtle yet potent force that can shape the outcomes of investigations.
+Sunde and Dror (2021) explored the susceptibility of digital forensics examiners to bias, revealing how preconceived notions and contextual information can skew their observations and interpretations.
+Despite the digital evidence’s facade of objectivity, the human factor introduces variability, leading to inconsistent conclusions among experts analyzing identical datasets (Sunde and Dror, 2021).
 3.1 Confirmation Bias in Text and Face Recognition
+Figure 2: John asked Lucy for help with a license plate, reading “BC 537”, but unsure about the last two characters and the first one. Credit: Fontani (2021).
+Figure 3: Image processing could lead to produce a noticeable different image, with different face characteristics. Credit: Fontani (2021).
 Fontani’s blog post on Amped Software depicted a hypothetical situation featuring two characters, John and Lucy (Figure 2(Fontani, 2021).
 In this scenario, John inadvertently influenced Lucy’s interpretation of a license plate by prematurely sharing his own interpretation.
-This phenomenon, known as confirmation bias, may lead Lucy to unconsciously process pixels and select frames that align with John’s interpretation.
-The post further delved into the intricacies of face comparison, noting that varying processing techniques can result in significantly different facial appearances.
+The cognitive bias present in this example could be the confirmation bias, since this latter is a type of cognitive bias where individuals are more likely to seek out, interpret, and remember information that confirms their pre-existing belief (Nickerson, 1998; Cooper and Meterko, 2019; Meterko and Cooper, 2022).
+John’s premature sharing of his interpretation could lead Lucy to interpret the license plate in the same way, confirming John’s interpretation rather than considering other possible interpretations.
+It may lead Lucy to unconsciously process pixels and select frames that align with John’s interpretation.
+The post further exploited into the intricacies of face comparison, noting that varying processing techniques can result in significantly different facial appearances.
 It underscored the necessity of withholding the suspect’s face from the examiner prior to the enhancement process to prevent unconscious bias towards a match.
@@
 Therefore, it’s crucial that the examiner doesn’t see the suspect’s face before the enhancement process to avoid unconsciously adjusting the enhancement to create a match.
-This unconscious adjustment is a form of cognitive bias.
-Figure 2: John asked Lucy for help with a license plate, reading “BC 537”, but unsure about the last two characters and the first one. Credit: Fontani (2021).
-Figure 3: Image processing could lead to produce a noticeable different image, with different face characteristics. Credit: Fontani (2021).
 Furthermore, Sunde and Dror (2019) underscored the importance of digital forensics as a rapidly growing field within forensic science.
@@
 (2013).
-The enhancement of images extracted from video cameras can lead to a degradation of the overall quality of information.
-This is due to factors such as excessive compression, distance from the recording plane, and limited overall resolution.
-International best practices suggest verifying on a case-by-case basis whether the level of information is sufficient to extract useful data for investigations (Khin et al., 2020).
+The enhancement of images extracted from video cameras can indeed lead to a degradation of the overall quality of information.
+This degradation can be attributed to factors such as excessive compression (Maity et al., 2023), distance from the recording plane (Wang et al., 2023), and limited overall resolution (Maity et al., 2023; Wang et al., 2023).
+International best practices suggest verifying on a case-by-case basis whether the level of information is sufficient to extract useful data for investigations (Tenopir et al., 2020; Soltani and Nikou, 2020).
 However, when examining low-contrast images that present pseudo-random visual patterns after an initial enhancement process, it is crucial to mitigate the risk of pareidolia.
-This bias is particularly potent when the object of interest refers to “human faces” or more generally to “letters/numbers” or known human structures (Wang and Yang, 2018; Zhou and Meng, 2020a).
+This bias is particularly potent when the object of interest refers to “human faces” or more generally to “letters/numbers” or known human structures (Wang and Yang, 2018; Zhou and Meng, 2020).
 Pareidolia is a subconscious illusion that tends to associate random shapes with known forms, especially human figures and faces.
 Classic examples include seeing animals or human faces in clouds, or a human face on the moon.
-In forensic investigations, we also suggest to entrust the analysis and interpretation to automatic methods or experts who can follow a “blind testing” approach, i.e., an interpretation detached from the knowledge of details and the reference context.
+In forensic investigations, we also suggest to entrust the analysis and interpretation to automatic methods (Solanke and Biasiotti, 2022) or experts who can follow a “blind testing” approach, i.e., an interpretation detached from the knowledge of details and the reference context (Cowan and Koppl, 2011; Servick, 2015).
 This helps to ensure the search for “certain” evidence is as rigorous and unbiased as possible.
-Zhou and Meng (2020b) discussed how some individuals exhibit a looser decision criterion for detecting faces, making them more prone to perceive faces where none exist.
+Zhou and Meng (2020) discussed how some individuals exhibit a looser decision criterion for detecting faces, making them more prone to perceive faces where none exist.
 This relates to a concept in signal detection theory known as response bias (Nguyen and Beins, 2013).
@@
 Just as some are more likely to see faces in random patterns due to biases in how they set thresholds for face judgments, forensic analysts could have biases influencing how strictly they apply standards of evidence to digital artifacts.
-Understanding individual differences in cognitive biases like threshold placement could help address potential sources of error and increase objectivity in digital forensic examinations.
-Furthermore, we present a case study which involved the analysis of surveillance camera footage to determine the presence of a passenger in a vehicle involved in a legal case.
-The analysis included scrutiny of various cameras and raised concerns about the validity of conclusions drawn by experts.
-The frames are part of a detailed analysis of the vehicle’s passages from different surveillance cameras, considering image overlaps and real passage times, including sunset.
-The aim was to evaluate the methodologies and conclusions of technical consultants and expert witnesses regarding the analysis of vehicle images, highlighting the lack of scientific rigor and proposing new analyses suggesting the absence of a passenger in the vehicle.
-Specifically, Figure 6 shows a sequence of frames that have been processed by a technical consultant.
-The frames demonstrate a clear reflection effect that disappears as the vehicle moves.
-The effect is attributed to the ease with which contour lines can introduce aberrations.
-The frames also exhibit strong chromatic variability due to video compression, which is known to heavily penalize the color component.
+Understanding individual differences in cognitive biases like threshold placement could help address potential sources of error and increase objectivity in digital forensic examinations (Berthet, 2021; Horsman, 2024).
+3.3 Case Study: Confirmation Bias and Pareidolia in Surveillance Camera Footage
+We present a case study with the objective of determining the presence of a passenger in a vehicle involved in a murder criminal case, through the analysis of surveillance camera footage.
+The methodology included the scrutiny of various cameras and a detailed analysis of the vehicle’s passages from different surveillance cameras, considering image overlaps and real passage times, including sunset.
+The data for this study is primarily derived from the surveillance camera footage.
+Figure 6 shows the only sequence of frames (among all the collected surveillance footage) in which the passenger’s presence is doubtful, and that has been processed by a technical consultant.
+These frames demonstrate a clear reflection effect that disappears as the vehicle moves, and exhibit strong chromatic variability due to video compression.
+The technical advisor claimed to discern a human face in these images, a perception potentially induced by confirmation bias, as his intervention was required precisely to identify the possible presence of the passenger.
+Specifically, on the passage of the car identified by the technical advisor he identifies with certainty, in the first frames of the passage, a face as a “clear” silhouette traceable to a passenger placed in the right front seat.
+A clear spot can be seen that he traces back to a silhouette.
+This spot disappears completely in the second part of the passage and then reappears but in inverted colors, that is, it becomes dark.
+In reality, the images contain pseudo-random blobs, and their temporal persistence can be attributed to a simpler and more evident reflection.
+The study highlights the lack of scientific rigor in these approaches and proposes new analyses that suggest the absence of a passenger in the vehicle.
 Figure 6: Sequence of frames that have been processed by the technical consultant.
-The technical advisor claimed to see the image of human faces in such images, a factor induced by confirmation bias: in reality, what is present in the images are pseudo-random blobs and their temporal hold is instead to be traced to a simpler and more obvious reflection.
-To demonstrate this, we conducted an experiment with college students.
-In the experiment, students were presented with frames from surveillance camera footage and tasked with identifying and interpreting the images.
-Crucially, no prior information was provided to the students to prevent any preconceived notions or confirmation biases from influencing their observations.
-The primary focus was on the students’ ability to discern whether a vehicle contained a passenger or not.
-The experiment was designed to rigorously assess the students’ cognitive perception, comprehension of the images, and their capacity to make accurate identifications solely based on the visual data at hand.
-This approach underscores the importance of maintaining an unbiased perspective in observational tasks.
-Table 2 summarizes the outcome of this analysis process, which reports a number of responses of 165 (15 subjects multiplied by the number of images viewed) in which in no case was the ”certain” presence of a human face evidenced.
 Image Students’ Evaluation
 Nothing: 8 Other: 7
-Those who answered ”other” identified:
+Those who answered “other” identified:
 • Automobile elements
 Nothing: 9 Other: 6
-Those who answered ”other” identified:
+Those who answered “other” identified:
 • Automobile elements
 Nothing: 10 Other: 5
-Those who answered ”other” identified:
+Those who answered “other” identified:
 • Silhouette with sun reflection
@@
 Nothing: 9 Other: 6
-Those who answered ”other” identified:
+Those who answered “other” identified:
 • A car seat and a person
@@
 Nothing: 13 Other: 2
-Those who answered ”other” identified:
+Those who answered “other” identified:
 • Indistinct silhouettes behind glass
 Nothing: 9 Other: 6
-Those who answered ”other” identified:
+Those who answered “other” identified:
 • There are two human silhouettes
 • Silhouettes of hands
-Table 2: Table representing students’ evaluation of the various images cutted from Figure 6.
-4 Bias Mitigating Strategies
-Mitigating these cognitive biases could significantly enhance decision-making across society, promoting long-term human well-being.
-Extensive research has been conducted on whether and how these biases can be mitigated (Acciarini et al., 2021).
-Korteling et al.
-(2021) conducted a systematic review of the literature on the retention and transfer of bias mitigation interventions, which yielded 12 peer-reviewed studies that adequately studied retention over a period of at least 14 days or transfer to different tasks and contexts.
-Most of these studies indicated that game-based interventions were effective after the retention interval and were more effective than video interventions.
-One study found indications of transfer of bias mitigation training across contexts.
-4.1 Biases Mitigating Strategies in Forensics
-Dror (2013) focused on strategies to mitigate these biases.
-He proposed that recognizing the spectrum of biases, not only those that can arise from knowing irrelevant case information, but also biases that emerge from base rate regularities, working “backwards” from the suspect to the evidence, and from the working environment itself, can strengthen forensic science.
-He suggested measures such as a triage approach and case managers to enhance the quality and effectiveness of the work carried out by forensic examiners.
+Table 1: Table representing students’ evaluation of the various images extracted from Figure 6.
+An experiment was conducted involving college students.
+The students were anonymously and without context presented with carefully selected frames from the surveillance camera footage, focusing on the most controversial and doubtful portions.
+The images shown to the students were a representative subset of the entire set of images extracted and processed by the technical advisor (Figure6).
+To ensure impartiality in their observations, no preliminary information was provided to the students.
+As part of the experiment, a questionnaire was administered for each image, consisting of the question: “What do you observe in this image?”, with response options such as “Nothing” and “Other,” along with an open text field for further clarification.
+This design encouraged a spectrum of responses and allowed students to give concise yet descriptive explanations.
+A total of 15 subjects participated in evaluating the selected images, resulting in 165 responses, as summarized in Table 1.
+Notably, only two of the responses conveyed a ”certain” identification of a human face.
+This underscores the necessity of meticulous and unbiased analysis when interpreting surveillance footage.
+The primary objective of the experiment was to assess the students’ ability to detect the presence of a passenger in the vehicle.
+By presenting them with ambiguous visuals, the experiment aimed to understand their perceptual limitations and biases.
+The absence of explicit certainty measurement in the questionnaire was addressed by considering the students’ chosen responses as an indication of their certainty levels.
+This highlights the importance of rigorous and unbiased analysis in interpreting surveillance footage.
+The study emphasizes the need to maintain an objective perspective in observational tasks (Altmann, 1974) and underscores the value of scientific rigor in such analyses.
+4 Strategies for Mitigating Cognitive Bias
+Among other fields, software engineering is currently experiencing a significant gap in the area of cognitive bias mitigation techniques, with a notable lack of both practical strategies and theoretical foundations (Mohanani et al., 2017).
+However, other fields have seen success in this area.
+For instance, in social work, the use of a nomogram tool and an online training course has been shown to effectively mitigate cognitive bias, leading to improvements in the accuracy of clinical reasoning (Featherston et al., 2019).
+Despite these advancements, it’s important to note that there is currently insufficient evidence to suggest that cognitive bias mitigation interventions significantly improve decision-making in real-life situations (Korteling et al., 2021).
+Nevertheless, the potential benefits of countering cognitive biases are clear.
+In healthcare, for example, addressing cognitive biases in decision-making can help reduce low-value care and enhance the impact of campaigns aimed at reducing such care (Scott et al., 2017).
+In the realm of gaming, the MACBETH serious game has been found to effectively mitigate cognitive biases such as the fundamental attribution error (Miller and Lawson, 1989) and confirmation bias.
+The game’s effectiveness is further enhanced through explicit instruction and repetitive play, which serve to reinforce learning (Dunbar et al., 2014).
+4.1 Strategies for Mitigating Cognitive Bias in Forensic Science
+In the field of forensic science, cognitive biases can significantly impact the accuracy and impartiality of examiner decisions.
+Specifically, Dror (2013) focused on strategies to mitigate confirmation bias, contextual influences (Nakhaeizadeh et al., 2014), and base-rate regularities (Thakur et al., 2021).
+Dror proposed that recognizing the spectrum of biases, not only those that can arise from knowing irrelevant case information, but also biases that emerge from base rate regularities, working “backwards” from the suspect to the evidence, and from the working environment itself, can strengthen forensic science.
+To mitigate these effects, several strategies have been proposed.
+First, cognitive training programs raise awareness among examiners about potential biases.
+Second, blind verification procedures, where the second examiner is unaware of the first examiner’s decision, help minimize bias.
+Third, linear examination processes, starting with evidence analysis before considering the suspect, reduce contextual contamination.
+Fourth, a triage approach tailors procedures based on case complexity, ensuring that resources are allocated effectively.
+Lastly, cognitive profiles aid in selecting the best-suited individuals for forensic work, enhancing overall objectivity and performance.
 In the pursuit of objectivity and accuracy in forensic analysis, the Forensic Science Regulator of the United Kingdom Government has proposed and implemented several strategies designed to ensure that the analysis is not influenced by any form of bias (Forensic Science Regulator, 2020).
-One of these strategies is the adoption of a structured and validated approach for the analysis and evaluation of evidence.
-Two models that exemplify this approach are the ACE-V model and the CAI model.
-The ACE-V model, which stands for Analysis, Comparison, Evaluation, and Verification, is a widely accepted method for comparing fingerprints (Reznicek et al., 2010).
-On the other hand, the CAI model, which stands for Case Assessment and Interpretation, is a framework based on Bayesian thinking (Jackson, 2011).
-It provides clarity on the role of forensic scientists within the criminal justice process and allows for the design of effective, efficient, and robust examination strategies.
-Documentation, independent verification, and maintaining ethical standards are other key strategies in the forensic process.
-Documentation involves taking contemporary notes and recording observations, decisions, and changes, ensuring transparency and enabling a thorough review.
-Independent verification of critical results by a competent peer, unaware of the original examiner’s judgment, adds an extra layer of scrutiny for accuracy.
-Furthermore, forensic operators must uphold ethical standards such as honesty, integrity, objectivity, and impartiality, and avoid conflicts of interest and external pressures to ensure unbiased and reliable analysis (Forensic Science Regulator, 2020).
-Lastly, controlling the flow of information to the analyst is paramount.
-This is achieved using methods such as “sequential unmasking” or “linear unmasking”, which help to prevent unnecessary information from influencing the analyst’s judgment.
-In this regard, Dror and Kukucka (2021) introduced Linear Sequential Unmasking–Expanded (LSU-E), a methodology aimed at improving decision-making in forensic science by reducing noise and bias.
-The approach advocates for an initial analysis of raw data or evidence without any reference material or context, followed by a sequential consideration of other relevant information based on its objectivity, relevance, and potential for bias.
-LSU-E is applicable to all types of decisions within forensic science, not just comparative ones.
-Its main objective is to optimize the sequence of information exposure to maximize information utility and minimize cognitive and psychological influences.
-In addition, LSU-E provides guidelines for determining the optimal sequence for presenting task-relevant information and emphasizes the need for transparent documentation of the influence and role of various pieces of information on the decision-making process.
-Camilleri et al.
-(2019) proposed that while some laboratories have implemented bias mitigating strategies with varying impact on operational efficiency, there has been no systematic assessment of the risk posed by cognitive bias.
-They suggested using a risk management framework to assess the potential impact of bias on forensic interpretations across multiple disciplines.
-This approach proved useful in assessing the effectiveness of existing bias mitigating strategies and identifying the latent level of risk posed.
-Furthermore, the authors affirm that all forensic organizations should seek to implement bias limiting measures that are simple, cost-effective, and do not adversely impact efficiency.
-Recently, Douglass et al.
-(2023) provided an examination of cognitive biases in digital forensics, particularly focusing on the potential of video-recorded eyewitness identification procedures for assisting evaluators in assessing eyewitness accuracy.
-The authors conducted three studies involving a total of 1,630 participants.
-The results revealed that evaluators, on average, successfully differentiated accurate from inaccurate witnesses based on videos of identification procedures alone.
-However, this ability was compromised when extraneous incriminating evidence was also provided.
-The authors found that instructions highlighting the limitations of forensic evidence did not preserve evaluators’ ability to discern accuracy when extraneous incriminating case evidence was provided.
-Moreover, case information affected other judgments, such as perceptions of the witness’s view.
-4.2 Biases Mitigating Strategies in Digital Forensics
-Grubl and Lallie (2022) addressed one of the most significant challenges in digital forensics: the precise age estimation of victims in child sexual abuse and exploitation cases.
-Investigators often need to determine the age of victims by examining images and interpreting the sexual development stages and other human characteristics.
-However, this process can be negatively impacted by a large forensic backlog, cognitive bias, and the immense psychological stress that this work can entail.
-In response to these challenges, the authors evaluated existing facial image datasets and proposed a new dataset tailored to the needs of similar digital forensic research contributions.
-This dataset, which included 327 images of individuals aged 0 to 20, is tested on the Deep EXpectation algorithm pre-trained on the IMDB-WIKI dataset (Grubl and Lallie, 2022).
-The results for young adolescents aged 10 to 15 and older adolescents/adults aged 16 to 20 are very encouraging, achieving Mean Absolute Errors as low as 1.79.
-Karie et al.
-(2019) also delved into the role of cognitive biases in digital forensics, particularly in the context of combating cybercrimes.
-The authors acknowledged that the increasing complexity of cyber-attacks and the vast amounts of data, also known as Big Data, that investigators need to sift through to unveil Potential Digital Evidence have made the process of forensic investigation more challenging.
-In response to these challenges, the authors proposed a generic framework for diverging DL cognitive computing techniques into Cyber Forensics (CF), referred to as the DLCF Framework.
-Such solutions can range from reducing bias in forensic investigations to challenging what evidence is considered admissible in a court of law or any civil hearing.
-5 The Art and Science of Deepfakes
+  1. 1.
+    Blinding Precautions:
+    1. (a)
+      Analysts should be shielded from information that is not directly relevant to the analysis.
+    2. (b)
+      Sequential unmasking can be used, where decisions on suitability are made before comparison with reference samples.
+    3. (c)
+      Careful records should be kept to ensure the order of disclosure and analysis is transparent.
+  2. 2.
+    Structured Approach (ACE-V and CAI):
+    1. (a)
+      ACE-V (Analysis, Comparison, Evaluation, and Verification) provides a structured process for fingerprint comparison (Reznicek et al., 2010).
+    2. (b)
+      CAI (Case Assessment and Interpretation) uses Bayesian thinking and balances prosecution and defense hypotheses (Jackson, 2011).
+    3. (c)
+      Both approaches emphasize transparency and avoid post hoc rationalization.
+  3. 3.
+    Awareness, Training, and Competence Assessment:
+    1. (a)
+      Practitioners need training on cognitive bias risks and mitigation strategies.
+    2. (b)
+      Proficiency testing and regular assessment help maintain competence.
+  4. 4.
+    Avoidance of Reconstructive Effects:
+    1. (a)
+      Contemporaneous notes or technical records prevent reconstructive bias.
+    2. (b)
+      Analysts should rely on memory as little as possible.
+  5. 5.
+    Avoidance of Role Effects:
+    1. (a)
+      Organizational structures should insulate scientists from potential biasing pressures.
+    2. (b)
+      Scientists must prioritize their duty to the court over any other obligations.
+Controlling the flow of information to analysts is crucial to prevent unnecessary influences on their judgment.
+Dror and Kukucka (2021) introduced Linear Sequential Unmasking–Expanded (LSU-E), a methodology that reduces noise and bias in forensic decision-making.
+LSU-E involves initial analysis of raw data without reference material, followed by a sequential consideration of relevant information based on objectivity and relevance.
+This approach optimizes information presentation to enhance utility and minimize cognitive biases.
+LSU-E also offers guidelines for documenting the influence of information on the decision-making process, ensuring transparency and accountability.
+Furthermore, Camilleri et al.
+(2019) emphasized the need for a systematic assessment of cognitive bias risks in forensic laboratories, proposing a risk management framework.
+Key mitigation strategies include raising awareness through training, developing guidance documents, and limiting access to task-irrelevant information.
+Redacting irrelevant details, implementing blind known tests, and conducting independent casefile reviews enhance objectivity and reduce expectation bias.
+These measures may ensure the integrity of forensic interpretations and minimize the impact of cognitive biases.
+5 The Intersection of Generative AI and the Craftsmanship of Deepfakes
+Generative Artificial Intelligence (GenAI) is an increasingly popular technology that has significant implications across various fields (Bockting et al., 2023; Stokel-Walker and Van Noorden, 2023).
+It refers to AI systems that can generate new content, such as text, images, and audio, in response to human prompts.
+These systems, including deepfakes and AI chatbots like Generative Pre-trained Transformers (e.g., GPT-4), use complex algorithms to produce outputs that are often indistinguishable from content created by humans (Stokel-Walker and Van Noorden, 2023).
+The technology is advancing rapidly, with each new version adding capabilities that increasingly encroach on human skills; however, the use of these “black box” AI tools can introduce biases and inaccuracies, potentially distorting scientific facts while still sounding authoritative.
 Figure 7: Statistics of papers published in the deepfake field.
@@
 (b) Numbers of papers published in: Article, Preprint, Proceedings, Chapter and Edited Book.
-The advent of advanced AI technologies has introduced new challenges in this field.
+Figure 8: A standard GAN framework.
+A Generator (G) creates data samples from noise, aiming to mimic the training set.
+A Discriminator (D) differentiates between real and G-generated data.
+Training ends when D can’t distinguish G’s images from training samples.
+The emergence of these sophisticated AI technologies has brought about fresh challenges in this domain.
 One such challenge is the detection of deepfakes, the research area of which is constantly expanding (as shown in Figure 7).
-Deepfakes are synthetic media created through generative models based mainly on Generative Adversarial Networks (GANs) and Diffusion Models (DMs) (Goodfellow et al., 2014; Sohl-Dickstein et al., 2015).
+Deepfakes are synthetic media created through generative models based mainly on Generative Adversarial Networks (GANs) and Diffusion Models (DMs) (Goodfellow et al., 2014).
 GANs are composed of a Generator (G) and a Discriminator (D) trained simultaneously through a competitive process.
@@
 When G creates images with the same data distribution as Ts, D will no longer be able to solve its task and the training phase can be considered completed.
-Currently, researchers demonstrated that synthetic images created by DMs (Sohl-Dickstein et al., 2015) are better than those generated by GAN engines in terms of photorealism, as the creation process follows a more accurate and “controlled” flow.
+Currently, researchers demonstrated that synthetic images created by DMs are better than those generated by GAN engines in terms of photorealism, as the creation process follows a more accurate and “controlled” flow.
 The basic idea of DMs is to iteratively add noise to an input random noise vector for synthetic data generation in order to model complex data distributions.
 Figure 8 and Figure 9 show generic GAN and DM schemes related to the creation of synthetic people’s faces.
-Deepfakes can pose significant challenges in distinguishing real images from manipulated ones, thereby complicating the task of digital forensic investigators: in fact, the problem of deepfake detection has been addressed extensively by the scientific community (Masood et al., 2023; Verdoliva, 2020).
+Figure 9: A Diffusion Model architecture.
+Training data is corrupted with added Gaussian noise.
+From this data ($x_{T}$ step), a reverse process is constructed to generate new samples resembling the original ones.
+Deepfakes can pose significant challenges in distinguishing real images from manipulated ones, thereby complicating the task of digital forensic investigators: in fact, the problem of deepfake detection has been addressed extensively by the scientific community (Masood et al., 2023; Verdoliva, 2020; Lin et al., 2024).
 In this context, preventing cognitive biases in digital forensics becomes crucial to ensure the objectivity and neutrality of judgments.
-Figure 8: A standard GAN framework including a Generator (G) that creates data samples from random noise, aiming to mimic the training set’s data distribution; a Discriminator (D) differentiates between real and G-generated data.
-Both are trained competitively.
-Figure 9: A generic Diffusion Model learning a latent variable model that maps the latent space with a fixed Markov chain.
-The data is corrupted by adding Gaussian noise gradually to get the approximate posterior $q(x_{t}|x_{t-1})$.
-The aim of training is to learn the inverse process $(p_{\theta}(x_{t-1}|x_{t}))$ so that new data can be generated by running the chain backward.
-The latent variables $x_{1}$, …, $x_{t}$ have the same dimension as the data $x_{0}$.
 6 The Impostor Bias: How AI Media Triggers Bias and Doubt in Perception
-The emergence of Generative AI (GenAI) has brought about a sea change in the multimedia landscape, opening up novel avenues for crafting and modifying content.
+The emergence of GenAI has brought about a sea change in the multimedia landscape, opening up novel avenues for crafting and modifying content.
 It is largely attributed to the development of a class of machine learning models known as foundation models (FMs) (Rabowsky, 2023).
 These models, which include the likes of ChatGPT released in November 2022, marked the beginning of a new era in Artificial Intelligence.
-Foundation models are distinguished by their powerful applications to GenAI, which involves the use of models to generate new content and transform existing content (Rabowsky, 2023).
+Foundation models are distinguished by their powerful applications to GenAI, which involves the use of models to generate new content and transform existing content.
 GenAI models can produce high-quality artistic media for visual arts, concept art, music, fiction, literature, video, and animation; distinguishing the real from the fake is becoming increasingly complex, as in the case of human face recognition and its veracity (Figure 10).
-Figure 10: Do these people exist? Credit: Guarnera et al. (2022)
+Figure 10: Figures 8 and 9 illustrate the architectures of generative models that produce these images.
+The verisimilitude of these faces could potentially lead observers to question the existence of the depicted individuals, thereby giving rise to the Impostor Bias.
+Credit: Guarnera et al.
+(2022)
 The existence of GenAI and its knowledge can lead to the development of a new type of cognitive bias, which we have identified as the “Impostor Bias”.
-This refers to an a priori distrust regarding the veracity of multimedia elements such as videos, photos, and audio, due to the awareness that these can be realistically generated by Artificial Intelligence models.
-Being subject to this bias, a condition towards which we believe we are increasingly heading in parallel with the increase in realism of GenAI multimedia products, entails and will entail a constant doubt that will also be influential when the victims are, for example, court operators, investigators, or magistrates themselves.
-In this regard, the technologies previously analyzed regarding deepfake detection will increasingly have to orient themselves not only towards the recognition of manipulated or altered multimedia artifacts, but also those generated from scratch starting from descriptive textual inputs.
-This is a great challenge on which we will have to work to avoid falling into cognitive traps dictated by prejudices of technological suspicion, risking declaring or considering original multimedia products instead generated by AI, and vice versa.
-Another risk stemming from GenAI pertains to the counterfeiting of artworks, the sale of forged masterpieces, and the infringement of copyright laws.
+This is a hypothetical bias with no empirical basis yet, and the term “Impostor” is derived from the “Impostor Syndrome”, a psychological phenomenon in which people doubt their competence and fear being exposed as fraudulent.
+This bias, however, refers to a different context: distrust of multimedia content generated by Artificial Intelligence.
+In the context of AI, “Impostor Bias” refers to the tendency to doubt the veracity of multimedia elements such as videos, photos and audios, due to the knowledge that these can be realistically generated by AI models.
+This bias manifests itself as an a priori distrust, regardless of the quality or context of the multimedia content.
+The name “Impostor Bias” is appropriate because, just as in Impostor Syndrome, there is a persistent doubt about the veracity of something - in this case, AI-generated media content.
+Although they may appear authentic and realistic, awareness of the possibility that they are AI-generated “impostors” can lead to doubt and distrust.
+The term “bias” is used to describe Impostor Bias because it refers to a systematic and predictable tendency in the way people perceive media content, regardless of the evidence presented.
+For instance, this is not simply a variation in evaluations due to different levels of knowledge of the evaluator.
+In other words, even when presented with AI-generated media content that is indistinguishable from the real thing, people with Impostor Bias may still doubt its authenticity due to the knowledge that such content may be AI-generated.
+This is a bias because it is based on an a priori assumption rather than an objective assessment of the content itself.
+Furthermore, the term “bias” implies that this tendency can lead to distortions in perception and judgement.
+For example, “Impostor Bias” may lead people to discard or devalue authentic media content because they perceive it as potentially false or misleading.
+Figure 11: Real vs deepfake images of famous artists.
+Judicial practitioners and investigators may struggle with determining the authenticity of multimedia content, especially as AI-generated media becomes more realistic and descriptive.
+The risk of falling into cognitive traps or suspecting original content as AI-generated underscores the complexity of this issue.
+Additionally, GenAI raises concerns about artwork counterfeiting, copyright infringement, and the potential for forged masterpieces to be sold as genuine.
 Indeed, such technologies are now capable of simulating the technique and artistic style of the most famous artists, thereby compromising the ability to correctly discern between real and simulated works (Epstein et al., 2023; Leotta et al., 2023) (Figure 11).
 Forensic examiners from 21 countries showed limited understanding and appreciation of cognitive bias, with fewer than half supporting blind testing, highlighting the need for procedural reforms to blind them to potentially biasing information (Kukucka et al., 2017).
-Figure 11: Real vs deepfake images of famous artists.
-7 Deepfake Detection Methods
+7 Generative AI and Deepfake Detection Methods
+The mitigation of the potential impacts of the hypothetical “Impostor Bias” could significantly rely on Generative AI and deepfake detection methods.
+As AI-generated media products become increasingly realistic (Verdoliva, 2020; de Lima-Santos and Ceron, 2021), the prevalence of “Impostor Bias” is expected to rise, posing significant challenges in various sectors, including the justice system.
+In this context, technologies for deepfake detection become increasingly important.
+These technologies are evolving to not only recognize manipulated or altered multimedia artifacts but also those generated from scratch using descriptive textual inputs.
+Detecting content generated by Generative AI, such as ChatGPT, is a rapidly evolving field of study.
+Recent research has focused on developing machine learning tools capable of distinguishing between human-written text and machine-generated content (Prillaman, 2023; Nature, 2023; Weber-Wulff et al., 2023; Perkins et al., 2024; Liu et al., 2024).
+These tools, such as the one described in a study by Perkins et al.
+(2024), analyze various features of writing style, including variation in sentence lengths, and the frequency of certain words and punctuation marks.
+However, the efficacy of these detection tools can be significantly reduced when they are confronted with machine-generated content that has been modified using techniques designed to evade detection (Perkins et al., 2024).
+For instance, the study found that the detectors’ already low accuracy rates (39.5%) showed major reductions in accuracy (17.4%) when faced with manipulated content.
+Despite these challenges, the development of more accurate and robust GenAI detection methods continues to be a critical area of research, given the increasing prevalence of AI-generated content in various domains (Perkins et al., 2024; Liu et al., 2024).
+For example, a tool called CheckGPT has been developed, which examines 20 features of writing style to determine whether an academic scientist or ChatGPT wrote a piece of text (Liu et al., 2024).
+The tool was found to be highly accurate, achieving an average classification accuracy of 98% to 99% for task-specific discipline-specific detectors and the unified detectors.
 Figure 12: Fourier spectrum of different categories of data: (a) real images; (b) images generated by two GAN architectures (GDWCT (Cho et al., 2019) and StarGAN (Choi et al., 2018)); (c) images generated by two DM architectures (Latent Diffusion (Ramesh et al., 2022) and GLIDE (Nichol et al., 2022)).
 The abnormal frequencies (light peaks) are mainly visible in the images generated by artificial intelligence engines.
-Researchers have demonstrated that generative engines leave traces on synthetic content that can be identified and detected in the frequency domain (Guarnera et al., 2020c; Zhang et al., 2019; Marra et al., 2019; Giudice et al., 2021; Dzanic et al., 2020; Durall et al., 2020) (Figure 12).
-These traces are characterized by both the network architecture (number and type of layers) and its specific parameters (Yu et al., 2019a).
+Furthermore, researchers have demonstrated that generative engines leave traces on synthetic content that can be identified and detected in the frequency domain (Guarnera et al., 2020c; Zhang et al., 2019; Marra et al., 2019; Giudice et al., 2021; Dzanic et al., 2020; Durall et al., 2020) (Figure 12).
+These traces are characterized by both the network architecture (number and type of layers) and its specific parameters (Yu et al., 2019).
 In order to distinguish real data from deepfake, Guarnera et al.
@@
 Vision Transformer-based solutions for deepfake detection have also been proposed (Wodajo and Atnafu, 2021; Coccomini et al., 2022; Heo et al., 2023; Wang et al., 2022)Wodajo and Atnafu (2021) combined transformers with a convolutional network to extract patches from detected faces in videos.
-Several studies (Yu et al., 2019b, 2021; Girish et al., 2021; Asnani et al., 2023; Yu et al., 2020; Guarnera et al., 2022) have explored identifying specific GAN models used in creation (Model Attribution Task).
+Several studies (Yu et al., 2019, 2021; Girish et al., 2021; Asnani et al., 2023; Yu et al., 2020; Guarnera et al., 2022, 2024) have explored identifying specific GAN models used in creation (Model Attribution Task).
 Guarnera et al.
@@
 Sha et al. (2023) GLIDE, Latent Diffusion, Stable Diffusion, DALL·E 2 MSCOCO (a), Flickr30k (b) 90.2%(a), 84.6% (b)
-Table 3: A summary of the discussed deepfake detection methods. Legend: AUC = Area Under the Curve; TAR = Transfer learning-based Autoencoder with Residuals; ViT = Vision Transformer.
+Table 2: A summary of the discussed deepfake detection methods.
 Experimental results of all these methods show that, in general, all generative models leave unique traces that can solve all previously listed tasks with high accuracy.
@@
 However, we want to highlight an important element of the previously listed methods and not just deepfake detection.
-All methods in the literature achieve extremely high results in ”constrained” contexts, that is, with architectures known a priori.
+All methods in the literature achieve extremely high results in “constrained” contexts, that is, with architectures known a priori.
 In practice, current deepfake detectors fail to generalize with synthetic images generated by novel architectures (different from those used during the training procedure), resulting in a drastic drop in classification performance.
@@
 8 Discussion
-We highlighted the importance of deepfakes in the field of digital forensics, which we consider to be a topic of great relevance and timeliness: in particular, the concept of impostor bias that we introduced is increasingly prevalent, although it had not yet been named and defined precisely.
-A concrete example of impostor bias was seen with the recent war in Ukraine, the first war characterized by constant propaganda from both sides carried out on all media, and especially on social networks (Ciuriak, 2022; Suciu, 2022).
-Being flooded with videos, photos and statements inevitably leads to believe a priori that everything that is seen can be modified, or generated by AI models through a simple prompt (Linehan et al., 2023).
-The habit of suspicion developed during this “digital warfare” with propaganda and false information led to immediate doubt of all published media, sometimes leading to the discovery that some of this media was deepfake (Bond, 2023).
-This condition of prior doubt on the veracity of the media will be more and more invasive and constant, especially when it comes to delicate material, such as a communication from a head of state in wartime (Linehan et al., 2023): for this reason, we believe that it is necessary to use the best tools for detecting deepfakes and AI-generated multimedia to avoid being deceived, especially to avoid considering fake multimedia that, paradoxically, are instead real.
-Therefore, the data presented herein, along with previous findings, underscore the imperative need for further exploration of cognitive biases in digital forensic science.
-The focus should not be solely on the biases themselves, but rather on the strategies that can be employed to counteract and disengage from them.
-One potential approach could involve the creation and implementation of specialized training programs for forensic practitioners.
-These programs would aim to address cognitive biases in various scenarios, such as those examined in this review.
-By enhancing awareness of cognitive biases, we could potentially reduce the likelihood of practitioners being influenced or overwhelmed by these biases.
-This would enable them to disregard irrelevant elements of the investigation, separate personal experiences and other contextual factors that may “contaminate” decision-making processes, thereby preserving the objectivity of their judgments.
+The concept of Impostor Bias, though not yet widely recognized, is increasingly relevant in the era of deepfakes and digital forensics.
+The recent Ukraine war, marked by propaganda and misinformation on social media (Ciuriak, 2022; Suciu, 2022), has heightened this bias.
+The constant exposure to videos, photos, and statements fosters an inherent suspicion of media manipulation or AI generation (Linehan et al., 2023).
+This ”digital warfare” has led to immediate doubt and scrutiny of all published media, sometimes uncovering deepfakes (Bond, 2023).
+The prevalence of Impostor Bias will persist, especially with sensitive content like wartime communications (Linehan et al., 2023).
+Thus, novel techniques and strategies are essential for effective mitigation.
+The unique nature of the Impostor Bias necessitates specialized approaches beyond the strategies discussed in Section 4.
+While those strategies offer valuable insights, addressing Impostor Bias effectively requires a combination of targeted measures.
+This includes shielding analysts from irrelevant information and providing specific training on cognitive bias risks.
+Additionally, proficiency testing and regular assessments maintain practitioner competence.
+However, to truly counter Impostor Bias, advanced technical tools for synthetic content detection are essential.
+Practical deep learning algorithms can detect generative models’ traces, aiding investigators in distinguishing authentic evidence.
+Given the success of deepfake detection algorithms (Pei et al., 2024; Gong and Li, 2024), these tools can assist in counteracting Impostor Bias and other digital biases.
+Similar detection methods for forgery images (Baumy et al., 2022; Zanardelli et al., 2023; Singh and Kumar, 2024) and multimedia data manipulation (Galante et al., 2023; Dunsin et al., 2024) further reinforce the value of algorithmic assistance in avoiding bias.
+The increasing sophistication of digital data demands the use of advanced deepfake detection techniques to maintain trust and accuracy in forensic examinations.
+Detection methods play a crucial role in identifying deepfakes, but interpreting and explaining the results is equally important.
+While some methods use classic machine learning techniques, others employ deep learning techniques, each with its advantages and limitations.
+Classic machine learning approaches are more interpretable but lack robustness, while deep learning methods offer higher performance and robustness but are more complex to explain.
+The question of trusting the predictions of detection algorithms is indeed a separate but related study, as explored by Mathews et al.
+(2023), Lim et al.
+(2022), and others (Cantero-Arjona and Sánchez-Macián, 2024; Pinhasov et al., 2024; Pandey et al., 2024; Pontorno et al., 2024).
+These studies enhance the “trust” in detection methods and underscore the need for further exploration of cognitive biases in digital forensic science.
+Creating specialized training programs for practitioners can enhance awareness, reduce the influence of biases, and preserve the objectivity of judgments by disregarding irrelevant elements and separating personal experiences.
 9 Conclusions and Future Works
-This article explores cognitive biases in forensics and digital forensics, including confirmation bias, expectation bias, and overconfidence in errors.
-It discusses methods to reduce these biases, such as game-based interventions and the Linear Sequential Unmasking-Expanded approach.
-The article highlights the challenge of deepfakes, synthetic media that can manipulate or impersonate individuals, threatening the integrity of digital evidence.
-It surveys current deepfake detection methods and their limitations.
-The article introduces the impostor bias, a new cognitive bias that may lead to false negatives and reduced confidence in digital forensic findings due to the prevalence of deepfakes.
-We conclude by proposing new research directions in digital forensics, suggesting the creation of advanced deepfake detection methods, exploring the factors contributing to impostor bias, and recommending interventions to mitigate this bias.
-We also underscore the need to examine the ethical, legal, and social implications of deepfakes.
+In this article, we have explored the significant impact of cognitive biases in the fields of forensics and digital forensics, specifically focusing on confirmation bias, anchoring bias, and hindsight bias.
+To reduce the influence of these biases, we have discussed strategies such as game-based interventions and the Linear Sequential Unmasking-Expanded approach.
+With the growing prevalence of deepfakes, synthetic media that can manipulate or impersonate individuals, the integrity of digital evidence is at risk.
+We have surveyed current deepfake detection methods and their limitations, introducing the novel concept of Impostor Bias.
+This bias, influenced by the widespread use of deepfakes, may lead to false negatives and reduced confidence in digital forensic findings.
+To address these challenges, we propose future research directions, including the development of advanced deepfake detection methods and a deeper understanding of the factors contributing to Impostor Bias.
+Furthermore, we emphasize the importance of interventions to mitigate this bias and the need to explore the ethical, legal, and social implications of deepfakes.
+By doing so, we aim to enhance the reliability and integrity of digital forensic practices, ensuring the accuracy and objectivity of forensic investigations.
 10 Acknowledgments